OpenStudy (unknownrandom):

Need some help on this. ^4sqrt((2/3))

OpenStudy (unknownrandom):

Here is a pic of the problem.

OpenStudy (anonymous):

\[\sqrt[4]{\frac{2}{3}}\] what are you supposed to do with it?

OpenStudy (unknownrandom):

Simplify

OpenStudy (anonymous):

there is no such mathematical operation as "simplify" perhaps you are supposed to get the fraction out of the radical? rationalize the denominator?

OpenStudy (unknownrandom):

Yes

OpenStudy (anonymous):

then multiply top and bottom by \(\sqrt[4]{3^3}=\sqrt[4]{27}\)

OpenStudy (anonymous):

that way the denominator will be \(\sqrt[4]{3^4}=3\)

OpenStudy (anonymous):

and the final answer will be \[\frac{\sqrt[4]{54}}{3}\] tell your math teacher this is "simplest radical form" or "rationalize the denominator" not "simplify"

OpenStudy (unknownrandom):

Thanks. I'm still confused on why you multiply by those numbers?

OpenStudy (anonymous):

you multiply by that number because that is the number of the denominator. when you multiply a radical by itself you no longer have a radical, you just have the number of the radical which is called rationalized